OpenStudy (anonymous):

A pancake recipe asks for three and one third times as much milk as flour. If one and one quarter cups of milk is used, what quantity of flour would then be needed, according to the recipe?

OpenStudy (anonymous):



OpenStudy (anonymous):

@ganeshie8 corrct?

ganeshie8 (ganeshie8):

wow ! thats almost correct very good work indeed :D

ganeshie8 (ganeshie8):

let me show where u did mistake

ganeshie8 (ganeshie8):

three and one third times as much milk as flour \(\large \frac{3\frac{1}{3}}{1}\)

ganeshie8 (ganeshie8):

one and one quarter cups of milk is used \(\large \frac{1\frac{1}{4}}{x}\)

ganeshie8 (ganeshie8):

set both ratios equal

ganeshie8 (ganeshie8):

\(\large \frac{3\frac{1}{3}}{1} = \frac{1\frac{1}{4}}{x}\)

ganeshie8 (ganeshie8):

\(\large x = \frac{1\frac{1}{4}}{3\frac{1}{3}}\)

ganeshie8 (ganeshie8):

\(\large x = \frac{\frac{5}{4}}{\frac{10}{3}} \)

ganeshie8 (ganeshie8):

\(\large x = \frac{5}{4} \times \frac{3}{10} \)

ganeshie8 (ganeshie8):

\(\large x = \frac{3}{8} \)

ganeshie8 (ganeshie8):

so three eigths of flour is required

ganeshie8 (ganeshie8):